If the measure of the congruent legs of an isosceles triangle is doubled, the perimeter will be 36 cm. If each of the congruent legs of the triangle is increased by 5 cm, the perimeter will be 32 cm. Find the length of the side of a triangle.
Answer by mananth(16946)

let the congruent legs be x each



third side be y
2x+2x+y = 36
4x+y=36....................1
increased by 5
x+5+x+5+y =32
2x+y+10=32
2x+y=22.................2
subtract (2) from (1)
2x=36-22
2x=14
x= 7
plug value of x in (1)
28+y=36
y=8
congruent sides = 7cm third side = 8cm
m.ananth@hotmail.ca